OpenStudy (anonymous):

What function is equal to cos(72°)?

OpenStudy (joannablackwelder):

sin(12 deg)

OpenStudy (anonymous):

\[\text{Cos}[72{}^{\circ}]=\sqrt{1-\text{Sin}[ 72{}^{\circ}]^2} \]\[\frac{1}{4} \left(-1+\sqrt{5}\right)=\sqrt{\frac{3}{8}-\frac{\sqrt{5}}{8}} \]\[\left\{\frac{1}{4} \left(-1+\sqrt{5}\right),\sqrt{\frac{3}{8}-\frac{\sqrt{5}}{8}}\right\}\text{//}\text{FullSimplify} \]\[\left\{\frac{1}{4} \left(-1+\sqrt{5}\right),\frac{1}{4} \left(-1+\sqrt{5}\right)\right\} \]\[\{0.309017,0.309017\} \]